Aqual:

2x+2y=5 3x+3y=7 How many solutions?

Mercury:

old question, but to help out future users: easiest way would be to graph and just see how many times the lines intersect (either 0, 1, or infinitely many) but it would also be good to learn how to do this algebraically 2x+2y=5 3x+3y=7 you can multiply the first equation by 3 and the second equation by 2 to get coeffiicents to be 6 6x + 6y = 15 6x + 6y = 14 subtract the equations, then it'll be clear how many solutions there are